# Dequantified Diffusion-Schrödinger Bridge for Density Ratio Estimation

### v2

@@ 3.2 Gaussian Dequantization for Mollified Time Score
Let $\{\mathbf{X}_{t}\}_{t\in[0,1]}$ be a DI defined in Equation 8. Under A.1 and A.2, the time score for any $t\in(0,1)$ satisfies:
@@
where $L$ is the Lipschitz constant in A.2. Moreover, if $L<1$, this lower bound diverges to infinity:
@@
Dequantified Diffusion Bridge Interpolant. To stabilize the time score near the boundary, Gaussian dequantization (GD) is introduced by adding controlled perturbations to boundary samples, effectively mollifying the boundary densities and resulting uniformly bounded time score across $[0,1]$ (see Corollary 3.8 for details).
Specifically, for $\mathbf{x}_{i}\sim q_{i}$, its dequantified form, $\mathbf{x}_{i}^{\prime}$, can be obtained by:
@@
where $\varepsilon\in\mathbb{R}_{+}$ is small. The resulting dequantified densities are obtained via Gaussian convolution: $q_{i}^{\prime}(\mathbf{x})=(q_{i}\ast\mathcal{N}(\mathbf{0},\varepsilon\mathbf{E}_{d}))(\mathbf{x})$. This smoothing ensures bounded time scores near $t=0$ and $t=1$ (see Theorem 3.7 and Corollary 3.8 for details), improving stability in DRE.
Incorporating GD into the DBI yields the dequantified diffusion bridge interpolant (DDBI), formulated as:
@@
The DDBI can be expressed in terms of the original DBI by defining perturbed variables as $\mathbf{X}_{i}^{\prime}=\mathbf{X}_{i}+\mathbf{Z}_{\varepsilon}$, where $\mathbf{Z}_{\varepsilon}\sim\mathcal{N}(\mathbf{0},\varepsilon\mathbf{E}_{d})$. This results in:
@@ 3.3 Optimal Transport Rearrangement
To further reduce estimation error and improve efficiency of DRE, we extend the probability path of DDBI by aligning it with the entropically regularized optimal transport (OT):
@@
Dequantified Schrödinger Bridge Interpolant. Applying OTR followed by DDBI with $\alpha_{t}=1-t$ and $\beta_{t}=t$ yields the dequantified Schrödinger bridge interpolant (DSBI):
@@
The probability path defined by DSBI solves the SB problem:
@@
Furthermore, to rigorously quantify the improvement brought by OTR, we establish the following result comparing the upper error bounds of DDBI and DSBI:
@@
Consider the DDBI and DSBI with coefficients $\alpha_{t}=1-t$, $\beta_{t}=t$. Let $\pi\in\Pi(q_{0}^{\prime},q_{1}^{\prime})$ be an arbitrary coupling for DDBI and $\pi_{2\gamma^{2}}$ the entropically OT coupling for DSBI as defined in XX. Define the error functional $\mathcal{E}:=\left|\int_{0}^{1}\partial_{t}\log q_{t}^{\prime}\,\mathrm{d}t-\log\frac{q_{1}^{\prime}}{q_{0}^{\prime}}\right|$. Then, the upper bound of $\mathbb{E}_{\pi_{2\gamma^{2}}}[\mathcal{E}_{\text{DSBI}}^{2}]$ is less than or equal to that of $\mathbb{E}_{\pi}[\mathcal{E}_{\text{DDBI}}^{2}]$.
@@ 3.4 Dequantified Diffusion Bridge DRE
Time Score-matching Loss. We train a time score model $s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)$ to approximate the time score $\partial_{t}\log q_{t}^{\prime}(\mathbf{x})$ by minimizing the time score-matching loss (Choi et al., 2022)
@@
However, $\partial_{t}\log q_{t}^{\prime}(\mathbf{x})$ is intractable in practice. To bypass this, an equivalent, integration-by-parts form (Song & Ermon, 2020; Choi et al., 2022) is proposed:
@@
Estimating Target Log Density Ratio. Given the optimal parameters $\boldsymbol{\theta}^{\star}$ obtained by minimizing $\mathcal{L}3$, the log density ratio at any point $\mathbf{x}$ can be estimated as:
$\log r^{\star}(\mathbf{x})\approx\int^{1}_{0}\partial_{t}\log q_{t}^{\prime}(\mathbf{x})\mathrm{d}t\approx\int^{1}_{0}s_{\boldsymbol{\theta}^{\star}}^{t}(\mathbf{x},t)\mathrm{d}t.$ (24)
@@
Algorithm 1 Pseudocode for training and estimation processes of $\text{D}^{3}\text{RE}$ with DDBI
@@
Algorithm 2 Pseudocode for training and estimation processes of $\text{D}^{3}\text{RE}$ with DSBI
@@ 4 Experiments
In experiments involving $\text{D}^{3}\text{RE}$, the DDBI and DSBI are implemented. Unless specified otherwise, we use the following settings: $\alpha_{t}=1-t,\beta_{t}=t,\gamma^{2}=0.5,\varepsilon=1e-5$ and $\lambda(t)=\gamma^{2}t(1-t)$. Under this configuration, the interpolant $\mathbf{I}(\mathbf{X}_{0},\mathbf{X}_{1},t)=(1-t)\mathbf{X}_{0}+t\mathbf{X}_{1}$ aligns with the Benamou-Brenier solution to the optimal transport problem in Euclidean space (McCann, 1997). The parameterized score model is trained with time score matching loss $\mathcal{L}_{3}$ and optimized with Adam optimization method.
@@ 4.1 Density Estimation
Energy-based Modeling on MNIST. We applied the proposed $\text{D}^{3}\text{RE}$ framework for density estimation on the MNIST dataset (LeCun, 1998), leveraging pre-trained energy-based models (EBMs) (Choi et al., 2022). Experimental details are in Section C.4. The results are reported in bits-per-dimension (BPD). Results in Tables 2 and 3 show that $\text{D}^{3}\text{RE}$ achieves the lowest BPD values across all noise types (Gaussian, Copula, and RQ-NSF), outperforming baselines and existing methods.
@@
Table 2: Comparison of the estimated densities on MNIST dataset based on pre-trained energy-based models. The results are reported in bits-per-dim (BPD). Lower is better. The reported results for NCE and TRE are sourced from (Rhodes et al., 2020).
@@ 4.2 Mutual Information Estimation
The evolution of estimated MI across epochs for $d=\{40,80,120\}$, comparing $\text{D}^{3}\text{RE}$ with DRE-$\infty$, are analyzed. Results in Figure 7 show that the red (DSBI) and green (DDBI) curves outperform the blue and yellow (DRE-$\infty$) curves in two aspects. First, $\text{D}^{3}\text{RE}$ converges to the true MI value more rapidly, as it expands trajectory sets (see Corollary 3.3), improving interpolation accuracy. Second, it exhibits greater stability with fewer fluctuations around the true MI, indicating more reliable estimates.
@@ 4.3 Analysis and Discussion
Ablation Study on $\gamma^{2}$. The ablation study on varying $\gamma^{2}$ values (Figure 8) shows distinct convergence behaviors in MI estimation. For all dimensions ($d=\{40,80,120\}$), smaller $\gamma^{2}$ values ($\leq 0.01$) lead to faster convergence, especially in lower dimensions ($d=40$), but excessively small values ($\gamma^{2}=0.001$) cause instability later. Larger $\gamma^{2}$ values ($\geq 0.1$) converge more slowly but stabilize over time, particularly in higher dimensions ($d=120$). $\gamma^{2}=0.1$ offers a balance between speed and stability across all dimensions, suggesting that moderate regularization ($\gamma^{2}=0.01,0.1$) provides the best MI estimation performance. Additional comparisons are provided in Section C.5.
@@
Ablation Study on OTR. We conduct an ablation study to evaluate the role of OTR, comparing models without OTR (baseline, DI), models with OTR (DI+OTR), and models from the $\text{D}^{3}\text{RE}$ framework (DDBI and DSBI). In Figures 2 and 5, DI generates distorted and misaligned intermediate distributions. This shows its limited ability to align with the target distribution. DI+OTR improves alignment but remains suboptimal. Models from the $\text{D}^{3}\text{RE}$ further enhance distribution quality, with DSBI achieving the most precise alignment. This underscores OTR’s crucial role in improving intermediate distributions.
@@
(a) $d=40$
(b) $d=80$
(c) $d=120$
Figure 7: Evolution of estimated MI across epochs with varying methods and dimensions $d=\{40,80,120\}$. $\text{D}^{3}\text{RE}$ outperforms the baseline in both speed and precision. The DRE-$\infty$ (Choi et al., 2022) is regarded as the ‘baseline’ in this experiment.
(a) $d=40$
(b) $d=80$
(c) $d=120$
@@ 5 Related Works
Density Ratio Estimation. DRE is an essential technique in machine learning but faces challenges in high-dimensional settings and when distributions are significantly different. Early methods (Sugiyama et al., 2012; Gutmann & Hyvärinen, 2012) often struggled with instability, known as the density-chasm problem, in these scenarios. To overcome these challenges, TRE (Rhodes et al., 2020), an extension of NCE (Gutmann & Hyvärinen, 2012), introduced a divide-and-conquer approach, breaking the problem into simpler subproblems for better performance. DRE-$\infty$ (Choi et al., 2022) further advanced this by interpolating between distributions through an infinite series of bridge distributions, improving stability and accuracy. F-DRE (Choi et al., 2021) used an invertible generative model to map distributions into a common feature space before estimation. Recent methods, such as (Kato & Teshima, 2021), have addressed overfitting in flexible models, and (Nagumo & Fujisawa, 2024; Luo et al., 2024) focused on improving robustness to outliers. MDRE (Srivastava et al., 2023) tackled distribution shift through multi-class classification, offering an alternative to binary classification in high-discrepancy cases. Additionally, geometric approaches, like (Kimura & Bondell, 2024), have enhanced DRE accuracy by incorporating the geometry of statistical manifolds. Building on these advancements, our work proposes a novel method to improve both the accuracy and robustness of high-dimensional DRE.
Diffusion Bridge. Denoising diffusion implicit models (DDIMs) (Song et al., 2020a) have been proposed as an efficient alternative to denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020), which require simulating a Markov chain for many steps to generate samples. Diffusive interpolants (Albergo et al., 2023) provide a unifying framework of flow-based models and diffusion models, bridging arbitrary distributions using continuous-time stochastic processes. DDBMs (Zhou et al., 2023) are proposed as a natural alternative to cumbersome methods like guidance or projected sampling in generative processes. Our proposed DDBI and DSBI build upon diffusive interpolants and DDBMs by incorporating Brownian bridge into the interpolation strategy construction.
@@ 6 Conclusions
In this work, we propose $\text{D}^{3}\text{RE}$, a unified, robust and efficient framework for density ratio estimation. It is the first framework to directly address both the density-chasm and support-chasm problems, enabling uniformly approximated density-ratio estimation (Proposition 3.5). By incorporating diffusion bridges and GD, we construct the DDBI, which expands support coverage (Theorem 3.2,Corollary 3.3) and stabilizes the time score near boundaries (Corollary 3.8). Building upon DDBI, the OTR is incorporated to derive the DSBI, which offers more efficient and accurate density ratio estimation (Theorem 3.7) by solving the Schrödinger bridge problem (Proposition 3.6). Together, DDBI and DSBI form the core of the $\text{D}^{3}\text{RE}$ framework, enabling uniformly approximated and stable density-ratio estimation (Proposition 3.5). Extensive experiments validate these findings, demonstrating the superior performance of $\text{D}^{3}\text{RE}$ in tasks like density-ratio estimation on synthetic data, mutual information estimation, and density estimation.
While $\text{D}^{3}\text{RE}$ advances the state of density-ratio estimation, several directions remain open. Future work may explore adaptive or learned solvers to reduce function evaluation overhead, as well as more expressive interpolants to further improve robustness in handling complex or multi-modal distributions.
@@ Acknowledgements
The work is supported by the Fundamental Research Program of Guangdong, China, under Grants 2023A1515011281.
@@ Appendix A Proofs
Based on Theorem 2.4 of (Léonard, 2014), (De Bortoli et al., 2021) established that the solution to the SB problem as detailed in Equation 19, $\pi^{\star}$, satisfies the SB conditions: (1) the optimization problem for $\pi^{\star}$ is equivalent to the entropically regularied OT problem, with the optimal coupling $\pi_{2\gamma^{2}}$ defined in Equation 17; (2) for samples $\left(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime}\right)\sim\pi^{\star}$, the associated conditional path distributions $\pi^{\star}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})$ minimize the KL divergence: $\mathbb{E}_{(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\sim\pi^{\star}}\mathrm{KL}\left(\pi^{\star}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\|\pi_{\mathrm{ref}}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\right)$, where $\pi_{\mathrm{ref}}$ is the reference path distribution satisfying $\log\pi_{\mathrm{ref}}(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})=\frac{\|\mathbf{x}_{0}^{\prime}-\mathbf{x}_{1}^{\prime}\|^{2}}{2\gamma^{2}}+\text{const}$. These conditional distributions are optimized using Brownian bridges of diffusion scale $\gamma$, conditioned on the endpoints $\mathbf{x}_{0}^{\prime}$ and $\mathbf{x}_{1}^{\prime}$. The marginal distribution at intermediate time $t$ along the Brownian bridge is given by $q_{t}(\mathbf{x}\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})=\mathcal{N}(\mathbf{x}\mid(1-t)\mathbf{x}_{0}^{\prime}+t\mathbf{x}_{1}^{\prime},t(1-t)\gamma^{2}\mathbf{E}_{d})$ (Tong et al., 2023). Since our proposed OTR method uses Sinkhorn’s algorithm to solve the entropically regularized OT problem, and the probability paths of our DDBI with $\alpha_{t}=1-t$ and $\beta_{t}=t$ align with those of the Brownian bridge, a trajectory $\mathbf{x}_{t}^{\prime}$ generated by first sampling $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\sim\pi^{\star}$, then sampling $\mathbf{x}_{t}\sim q_{t}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})$ satisfies the SB conditions, thus verifying the proposition. ∎

### v3

@@ 3.2 Gaussian Dequantization for Mollified Time Score
Let $\{\mathbf{X}_{t}\}_{t\in[0,1]}$ be a DI defined in Equation 8. Under A.1 and A.2, the time score for any $t\in(0,1)$ satisfies
@@
where $L$ is the Lipschitz constant in A.2. Moreover, if $L<1$, this lower bound diverges to infinity
@@
Dequantified Diffusion Bridge Interpolant. To stabilize the time score near the boundary, we introduce Gaussian dequantization (GD) by adding controlled perturbations to boundary samples, effectively handling the boundary densities and resulting uniformly bounded time score across $[0,1]$ (see Corollary 3.8 for details).
Specifically, for $\mathbf{x}_{i}\sim q_{i}$, its dequantified form $\mathbf{x}_{i}^{\prime}$ can be obtained by
@@
where $\varepsilon\in\mathbb{R}_{+}$ is small. The resulting dequantified densities are obtained via Gaussian convolution, $q_{i}^{\prime}=q_{i}\ast\mathcal{N}(\mathbf{0},\varepsilon\mathbf{E}_{d})$. This smoothing ensures bounded time scores near $t=0$ and $t=1$ (see Theorem 3.7 and Corollary 3.8 for details), improving stability in DRE.
Incorporating GD into the DBI yields the dequantified diffusion bridge interpolant (DDBI), formulated as
@@
The DDBI can be expressed in terms of the original DBI by defining perturbed variables as $\mathbf{X}_{i}^{\prime}=\mathbf{X}_{i}+\mathbf{Z}_{\varepsilon}$, where $\mathbf{Z}_{\varepsilon}\sim\mathcal{N}(\mathbf{0},\varepsilon\mathbf{E}_{d})$. This results in
@@ 3.3 Optimal Transport Rearrangement
To further reduce estimation error and improve efficiency of DRE, we extend the probability path of DDBI by aligning it with the entropically regularized optimal transport (OT)
@@
Dequantified Schrödinger Bridge Interpolant. Applying OTR followed by DDBI with $\alpha_{t}=1-t$ and $\beta_{t}=t$ yields the dequantified Schrödinger bridge interpolant (DSBI)
@@
The probability path defined by DSBI solves the SB problem
@@
Furthermore, to rigorously quantify the improvement brought by OTR, we establish the following result comparing the upper error bounds of DDBI and DSBI.
@@
Consider the DDBI and DSBI with coefficients $\alpha_{t}=1-t$, $\beta_{t}=t$. Let $\pi\in\Pi(q_{0}^{\prime},q_{1}^{\prime})$ be an arbitrary coupling for DDBI and $\pi_{2\gamma^{2}}$ the entropically OT coupling for DSBI. Define the error functional $\mathcal{E}:=\big{|}\int_{0}^{1}\partial_{t}\log q_{t}^{\prime}\,\mathrm{d}t-\log\frac{q_{1}^{\prime}}{q_{0}^{\prime}}\big{|}$. Then, the upper bound of $\mathbb{E}_{\pi_{2\gamma^{2}}}[\mathcal{E}_{\text{DSBI}}^{2}]$ is less than or equal to that of $\mathbb{E}_{\pi}[\mathcal{E}_{\text{DDBI}}^{2}]$.
@@ 3.4 Dequantified Diffusion Bridge DRE
Time Score-matching Loss. We train a time score model $s_{\boldsymbol{\theta}}^{t}(\mathbf{x},t)$ to approximate the time score $\partial_{t}\log q_{t}^{\prime}(\mathbf{x})$ by minimizing the time score-matching loss (Choi et al., 2022),
@@
However, $\partial_{t}\log q_{t}^{\prime}(\mathbf{x})$ is intractable in practice. To bypass this, an equivalent integration-by-parts form (Song & Ermon, 2020; Choi et al., 2022) is proposed
@@
Estimating Target Log Density Ratio. Given the optimal parameters $\boldsymbol{\theta}^{\star}$ obtained by minimizing $\mathcal{L}3$, the log density ratio at any point $\mathbf{x}$ can be estimated as
$\log r^{\star}(\mathbf{x})\approx\int^{1}_{0}\partial_{t}\log q_{t}^{\prime}(\mathbf{x})\mathrm{d}t\approx\int^{1}_{0}s_{\boldsymbol{\theta}^{\star}}^{t}(\mathbf{x},t)\mathrm{d}t,$ (24)
@@
Algorithm 1 Training and estimation of $\text{D}^{3}\text{RE}$ with DDBI
@@
Algorithm 2 Training and estimation of $\text{D}^{3}\text{RE}$ with DSBI
@@ 4 Experiments
For experiments involving $\text{D}^{3}\text{RE}$, we implement both DDBI and DSBI. Unless specified otherwise, we use the following settings: $\alpha_{t}=1-t,\beta_{t}=t,\gamma^{2}=0.5,\varepsilon=1e-5$ and $\lambda(t)=\gamma^{2}t(1-t)$. Under this configuration, the interpolant $\mathbf{I}(\mathbf{X}_{0},\mathbf{X}_{1},t)=(1-t)\mathbf{X}_{0}+t\mathbf{X}_{1}$ aligns with the Benamou-Brenier solution to the optimal transport problem in Euclidean space (McCann, 1997). The parameterized score model is trained with time score matching loss $\mathcal{L}_{3}$ and optimized with Adam optimization method.
@@ 4.1 Density Estimation
Energy-based Modeling on MNIST. We applied the proposed $\text{D}^{3}\text{RE}$ framework for density estimation on the MNIST dataset, leveraging pre-trained energy-based models (EBMs) (Choi et al., 2022). Experimental details are in Section C.4. The results are reported in bits-per-dimension (BPD). Results in Table 2 show that $\text{D}^{3}\text{RE}$ achieves the lowest BPD values across all noise types (Gaussian, Copula, and RQ-NSF), outperforming baselines and existing methods. See also Table 3 in the Appendix.
@@
Table 2: Comparison of the estimated densities on MNIST dataset based on pre-trained energy-based models. The results are reported in bits-per-dim (BPD). Lower is better. The reported results for NCE and TRE are from Rhodes et al. (2020).
@@
(a) $d=40$
(b) $d=80$
(c) $d=120$
Figure 3: Evolution of estimated MI across epochs with varying methods and dimensions $d=\{40,80,120\}$. $\text{D}^{3}\text{RE}$ outperforms the baseline in both speed and precision. DRE-$\infty$ (Choi et al., 2022) is regarded as the ‘baseline’ in this experiment.
@@ 4.2 Mutual Information Estimation
The evolution of estimated MI across epochs for $d=\{40,80,120\}$, comparing $\text{D}^{3}\text{RE}$ with DRE-$\infty$, are analyzed. Results in Figure 3 show that the red (DSBI) and green (DDBI) curves outperform the blue and yellow (DRE-$\infty$) curves in two aspects. First, $\text{D}^{3}\text{RE}$ converges to the true MI value more rapidly as it expands trajectory sets (see Corollary 3.3), improving interpolation accuracy. Second, it exhibits greater stability with fewer fluctuations around the true MI, indicating more reliable estimates.
@@ 4.3 Analysis and Discussion
[…]
@@
(a) $d=40$
(b) $d=80$
(c) $d=120$
The ablation study on varying $\gamma^{2}$ values (Figure 5) shows distinct convergence behaviors in MI estimation. For all dimensions ($d=\{40,80,120\}$), smaller $\gamma^{2}$ values ($\leq 0.01$) lead to faster convergence, especially in lower dimensions ($d=40$), but excessively small values ($\gamma^{2}=0.001$) cause instability later. Larger $\gamma^{2}$ values ($\geq 0.1$) converge more slowly but stabilize over time, particularly in higher dimensions ($d=120$). $\gamma^{2}=0.1$ offers a balance between speed and stability across all dimensions, suggesting that moderate regularization provides the best MI estimation performance. More results are provided in Section C.5.
@@
Ablation Study on OTR. We conduct an ablation study to evaluate the role of OTR, comparing models without OTR (baseline, DI), models with OTR (DI+OTR), and models from the $\text{D}^{3}\text{RE}$ framework (DDBI and DSBI). In Figures 2 and 7, DI generates distorted and misaligned intermediate distributions. This shows its limited ability to align with the target distribution. DI+OTR improves alignment but remains suboptimal. Models from the $\text{D}^{3}\text{RE}$ further enhance distribution quality, with DSBI achieving the most precise alignment. This underscores OTR’s crucial role in improving intermediate distributions.
@@ 5 Related Works
Density Ratio Estimation. DRE is an essential technique in machine learning but faces challenges in high-dimensional settings and when distributions are significantly different. Early methods (Sugiyama et al., 2012; Gutmann & Hyvärinen, 2012) often struggled with instability, known as the density-chasm problem, in these scenarios. To overcome these challenges, TRE (Rhodes et al., 2020), an extension of NCE (Gutmann & Hyvärinen, 2012), introduced a divide-and-conquer approach, breaking the problem into simpler subproblems for better performance. DRE-$\infty$ (Choi et al., 2022) further advanced this by interpolating between distributions through an infinite series of bridge distributions, improving stability and accuracy. F-DRE (Choi et al., 2021) used an invertible generative model to map distributions into a common feature space before estimation. Recent methods, such as Kato & Teshima (2021), have addressed overfitting in flexible models, and Nagumo & Fujisawa (2024); Luo et al. (2024) focused on improving robustness to outliers. MDRE (Srivastava et al., 2023) tackled distribution shift through multi-class classification, offering an alternative to binary classification in high-discrepancy cases. Additionally, geometric approaches, like Kimura & Bondell (2025), have enhanced DRE accuracy by incorporating the geometry of statistical manifolds. Building on these advancements, our work proposes a novel method to improve both the accuracy and robustness of high-dimensional DRE.
Diffusion Bridge. Denoising diffusion implicit models (DDIMs) (Song et al., 2020) have been proposed as an efficient alternative to denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020), which require simulating a Markov chain for many steps to generate samples. Diffusive interpolants (Albergo et al., 2023) provide a unifying framework of flow-based models and diffusion models, bridging arbitrary distributions using continuous-time stochastic processes. DDBMs (Zhou et al., 2024a) are proposed as a natural alternative to cumbersome methods like guidance or projected sampling in generative processes. Our proposed DDBI and DSBI build upon diffusive interpolants and DDBMs by incorporating Brownian bridge into the interpolation strategy construction.
@@ 6 Conclusions
In this work, we propose $\text{D}^{3}\text{RE}$, a unified, robust and efficient framework for density ratio estimation. It provides the first framework for directly addressing both the density-chasm and support-chasm problems, enabling uniformly approximated density-ratio estimation (Proposition 3.5). By incorporating diffusion bridges and GD, we construct DDBI, which expands support coverage (Theorem 3.2, Corollary 3.3) and stabilizes the time score near boundaries (Corollary 3.8). Building upon DDBI, OTR is incorporated to derive the DSBI, which offers more efficient and accurate density ratio estimation (Theorem 3.7) by solving the Schrödinger bridge problem (Proposition 3.6). Together, DDBI and DSBI form the core of the $\text{D}^{3}\text{RE}$ framework, enabling uniformly approximated and stable density-ratio estimation (Proposition 3.5). Extensive experiments validate these findings, demonstrating the superior performance of $\text{D}^{3}\text{RE}$ in tasks such as density-ratio estimation on synthetic data, mutual information estimation, and density estimation.
While $\text{D}^{3}\text{RE}$ advances density-ratio estimation methods, several directions remain open. Future work could explore adaptive or learned solvers to reduce function evaluation overhead, as well as more expressive interpolants to further improve robustness in handling complex or multi-modal distributions.
@@ Acknowledgements
This research work is supported by the Fundamental Research Program of Guangdong, China, under Grant 2023A1515011281.
@@ Appendix A Proofs
Based on Theorem 2.4 of (Léonard, 2014), (De Bortoli et al., 2021) established that the solution to the SB problem as detailed in Equation 19, $\pi^{\star}$, satisfies the SB conditions: (1) the optimization problem for $\pi^{\star}$ is equivalent to the entropically regularied OT problem, with the optimal coupling $\pi_{2\gamma^{2}}$ defined in Equation 17; (2) for samples $\left(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime}\right)\sim\pi^{\star}$, the associated conditional path distributions $\pi^{\star}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})$ minimize the KL divergence: $\mathbb{E}_{(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\sim\pi^{\star}}\mathrm{KL}\left(\pi^{\star}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\|\pi_{\mathrm{ref}}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\right)$, where $\pi_{\mathrm{ref}}$ is the reference path distribution satisfying $\log\pi_{\mathrm{ref}}(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})=\frac{\|\mathbf{x}_{0}^{\prime}-\mathbf{x}_{1}^{\prime}\|^{2}}{2\gamma^{2}}+\text{const}$. These conditional distributions are optimized using Brownian bridges of diffusion scale $\gamma$, conditioned on the endpoints $\mathbf{x}_{0}^{\prime}$ and $\mathbf{x}_{1}^{\prime}$. The marginal distribution at intermediate time $t$ along the Brownian bridge is given by $q_{t}(\mathbf{x}\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})=\mathcal{N}(\mathbf{x}\mid(1-t)\mathbf{x}_{0}^{\prime}+t\mathbf{x}_{1}^{\prime},t(1-t)\gamma^{2}\mathbf{E}_{d})$ (Tong et al., 2024). Since our proposed OTR method uses Sinkhorn’s algorithm to solve the entropically regularized OT problem, and the probability paths of our DDBI with $\alpha_{t}=1-t$ and $\beta_{t}=t$ align with those of the Brownian bridge, a trajectory $\mathbf{x}_{t}^{\prime}$ generated by first sampling $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\sim\pi^{\star}$, then sampling $\mathbf{x}_{t}\sim q_{t}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})$ satisfies the SB conditions, thus verifying the proposition. ∎